Twentieth Century Literature, vol. XLIII, 4, 1997, p. 394–409.

Edith Wharton and the Faubourg Saint-Germain: the diary of the Abbé Mugnier

Anne Foata

R.W.B. Lewis, in his 1975 groundbreaking biography of Edith Wharton,^{1} mentions the Abbé Mugnier, a curate at the Saint Thomas d’Aquin and Sainte Clotilde parishes in Paris, who was to become a lifelong friend of the novelist. Both churches are located in the heart of the Seventh Arrondissement on the Left Bank of the French capital which includes the Faubourg Saint-Germain section, the Parisian environment of the great hôtels particuliers (town mansions) of the French aristocracy. The “Faubourg” (as it is more familiarly known) is the setting of Henry James’s The American, where its haughty aloofness is epitomized by the old marquise de Bellegarde, and of parts of Wharton’s own The Custom of the Country during the heroine’s brief marriage to the count Raymond de Chelles.

The stronghold of the Legitimists (i.e., the nobility of the Ancien Régime), the “Faubourg” in due time had opened up to the newer nobility of the Napoleon and Louis-Philippe regimes; it had also become home to the wealthy industrialists of the mid and late nineteenth century and after the Franco-Prussian War to some of the more prominent writers and artists.^{2} Edith Wharton herself lived in its midst from 1906 to 1920, for short periods until 1911, and then on a more permanent basis after the sale of The Mount, her Lenox residence in Western Massachusetts.

Wharton was introduced to the life of its salons by the French writer and Academician Paul Bourget and his wife Minnie, the “Minnie-Pauls” as they appear in her letters and diaries until 1935, the year of Bourget’s death. Wharton’s social life in the Faubourg, which Lewis records with an eager relish for its aristocratic details, is quite staggering; of keener interest, however, for Wharton scholars is the number of her translators who were recruited amidst the Faubourg’s literati and nobility.

On Bourget’s suggestion, Charles du Bos, a friend of André Gide, translated The House of Mirth and later, Ethan Frome; Gide himself, who had professed the greatest admiration for the latter work, was approached by Wharton for Summer but eventually declined and du Bos did the translation (Lewis 398). The Vicomte Robert d’Humières, a friend of Marcel Proust, began work on The Custom of the Country, but was killed at the front without being able to finish it (Lewis 382). The eminent literary critic and art historian Louis Gillet on his own initiative offered to translate A Mother’s Recompense and later The Children, while one of his daughters, Louisette, busied herself with one of Wharton’s longer stories or novellas (Lewis 485). The Comtesse Jane d’Oilliamson, whom Wharton had first met when she was still the unmarried Jane de Polignac, translated “The Reckoning,” “The Confessional,” “Souls Belated” and other stories (Lewis 207 & 212). Charles du Bos, it is true, had an English mother and was bilingual, and so undoubtedly was Jane d’Oilliamson whose father, the Prince Alphonse, was born in England during his own father’s sojourn in this country as an émigré at the time of the French Revolution.^{∗}

All these names, with the exception of Jane d’Oilliamson and the addition of numerous others, many of which were to some degree or other known or familiar to Wharton, appear in the diary of the Abbé Mugnier, which was published in parts in 1985 under the title Le Journal de l’Abbé Mugnier (1879-1939).^{3} The Abbé, as he was known to the end of his life despite the fact that he had been made a Canon in 1925, knew about everybody.

He had been drawn into the milieu of the writers and artists in 1891 through his spiritual ministrations to the “decadent” author of A Rebours, J.K. Huysmans, whom he eventually converted to Roman Catholicism; as to the aristocrats, he started hearing the confessions of their footmen and parlor maids, went on with teaching the catechism to their sons and daughters, advised and comforted the wives, and ended up in their diningand drawing-rooms.

Not that he was quite without misgivings about his intense dining out. “No priest has ever dined more about town than I,” he confided to his diary on 29 January 1911, “I’m squandering my soul by platefuls.” Another time he ruefully admitted to himself that he was rather the “marriage at Cana” type of priest than the “fast in the desert” one (10 April 1912). He was an avid reader of the classics, a lifelong enthusiast of Chateaubriand and a great admirer of the music of Wagner; an untiring confidant of the gens du monde and gens de lettres, he did not, however, belong to the breed of the fashionable little society priests or court lackeys of eighteenth century boudoir fiction. With his rather shabby soutane and heavy brogues, his somewhat rustic appearance alone would have denied the fact, had not his mores been as unimpeachable as was his manner toward the intellectuals and aristos of his flock. His unpretentiousness was genuine; so were his curiosity for all things human and his benevolence. All his friends, high and low, knew they could tap his considerable knowledge of, and compassion for, the human heart.

The Abbé’s diary with its almost daily entries over sixty years spans what has been called la Belle Epoque, i.e., the late nineteenth century that came to an abrupt halt in 1914, and the troubled years of the entre-deux-guerres, the period between the two world wars. The Mercure de France edition presents a mere selection of this vast material, which the Abbé left to the first of his two “adopted nieces,” the nièce première as he called her, the Comtesse François de Castries. The nièce seconde or second niece by elective adoption, the Princesse Marthe Bibesco, published her own correspondence with the Abbé between 1951 and 1957,^{4} and a memoir Le Confesseur et les Poètes^{5} in 1970. A distant cousin of the poetess Anna de Noailles, a writer in her own right and noted socialite, the Princess who was acquainted with Walter Berry may well have met Edith Wharton; Wharton’s scholars might want to have a look at both works for possible references to their novelist.

The Abbé Mugnier died at ninety-one years of age, a mere four months before the Liberation of Paris in 1944; he had endured the hardships of no less than three major wars with one Revolutionary upheaval as well (the Commune de Paris of 1871).

In his biography of Edith Wharton, Lewis’s first mention of a joint attendance by both the Abbé and the novelist at a social event bears the date of 13 April 1908 and is recorded as a random example of Wharton’s social calendar for that day (Lewis 212). The Abbé Mugnier may have likewise made a note of the dinner at the Comtesse Rosa de Fitz-James’s, but if so the entry for that day or the next does not figure in the printed edition. Nor does Wharton’s name, by the way, in the entries of those early years of her sojourns in Paris. She appears for the first time in it in the entry for 21 April 1915 together with her beloved friend Walter Berry, and both of them are mentioned again in the entry for 13 August of the same year. There are three entries for 1918 (7 & 9 February, 22 March), where they figure together with that other dear friend of Wharton’s, Bernard Berenson. All three are mentioned again in the entries for October 1923 (Wharton and Berenson on the 11th, Wharton and Berry on the 16th). There is one last entry on 24 July 1924 with all three of them before Wharton’s name, and Berenson’s, disappear definitively. Walter Berry, who had been mentioned on his own on five occasions (25 June 1917, 17 April & 27 November 1918, 19 February & 17 October 1921) appears one last time on 11 June 1927, a few months before his death; we see him rescuing a bewildered Abbé from the private showing of Fritz Lang’s film Metropolis and driving him home to his rooms at the convent of the Sisters of Saint Joseph de Cluny on the rue Méchain.

There is a total of eight entries referring to Edith Wharton, eleven to Walter Berry and eight to Bernard Berenson, several of them, as we have seen, common to all two or three of them. These I will venture to translate from the French, for the benefit of those Wharton scholars who may be interested in catching an echo of Wharton’s “whirligig life,” as James is reported to have called the latter’s flurry of activity (Lewis 318). Numerous entries of the Abbé’s Journal refer at some length to Charles du Bos and Paul Bourget who were close friends of Wharton, in addition to writers such as André Gide, Jean Cocteau, Anna de Noailles, Paul Valéry and others who without being intimate were frequent visitors at Wharton’s various residences or co-guests at the houses of her friends. A generous listing of the members of the French aristocracy and of their salons which Wharton frequented can also be culled from it.

A notable silence on the Abbé’s part, however, and a missed opportunity on Wharton’s come to light in the pages of the diary, at least in its printed version. They concern Henry James with respect to the Abbé and Marcel Proust with respect to Wharton.

Of course, on the Abbé’s part it may well be that the entries recording an acquaintance with James were omitted in the printed edition of the Journal. We know how profusely James was wined, dined and teaed by all Wharton’s friends on his visits to Paris (Lewis 216 for instance). But then no entry of the diary mentions Wharton herself before 21 April 1915 and by that date the ailing James no longer came to Paris. Still, James’s absence from the Journal, except for the mere mention of his authorship of the biography of the American sculptor W.W. Story and a passing remark on his style, is a disappointment. The Abbé’s candid impressions of the rotund, magisterial, elderly James would have been gratifying.

There is no accountable reason for the missed opportunity du côté de chez Proust and Wharton. The former not only was a friend of the Abbé Mugnier but a close acquaintance of Walter Berry to whom he dedicated his Pastiches et Mélanges (Lewis 402); Wharton, moreover, might have met him in the salons of countless other mutual friends and acquaintances, Jean Cocteau’s for instance, the Charles du Bos’, Anna de Noailles’, to name only a few. That they shared the same world of princesses and duchesses need not be expatiated on, but that they might eventually have translated portions of each other’s work and failed to do so opens up an abyss of rueful frustration.

5 ANNE FOATA

Lewis mentions a letter from Gide to Wharton in March 1916 in which Gide suggests that Proust take over the translation of The Custom of the Country left unfinished by Robert d’Humières’ death, and actually gives her to understand that Proust is waiting for a word from her on that account (Lewis 400-401 & 557). More than a decade later, in 1930, Wharton was approached by Proust’s American publisher with the proposal that she assume the unfinished portion of Scott Moncrieff’s Remembrance of Things Past and translate Le Temps Retrouvé. Had she been “forty or fifty years younger,” as she confided in a letter to Berenson, she would have been gratified with the honor; at 68, however, she no longer felt herself quite able to muster the necessary energy (Lewis 494). C’est dommage!, as she might have said in the fluent French she used with many of her friends.

Here then are the references to Edith Wharton and her two close friends, Walter Berry and Bernard Berenson, as they appear in Le Journal de l’Abbé Mugnier. The ellipses in the text of the entries are my own, not the Abbé’s. They concern passages I deemed irrelevant

∗ I am indebted to Dr Sylvia Kahan of CUNY for information on Jane d’Oilliamson’s family origin.

References to Edith Wharton

21 April 1915

Had lunch at Mrs. Wharton’s^{6} with the Abbé Brémond^{7} and Berry. Mrs. Wharton was talking of the life of Story by Henry James. James’s work is like Péguy’s, she said, only much more complicated. Story’s life is quite extraordinary. He was an American sculptor, mediocre, who had established himself in Rome; a friend of the Brownings with whom he lived for a while. Has left no documents^{8} and yet James has written a remarkable story, a masterpiece.

13 August 1915

Dined last night at Walter Berry’s with Mrs. Wharton, Mr. & Mrs. Garrett, Rodier,^{9} Mrs. Astor, another American,^{10} and Bakst. On the table the severed heads of red roses swimming in a bowl of water, another array of charming heads! Berry’s apartment^{11} is the apartment of an amateur, a dilettante, with opened books scattered about, in every place and position, even on the floor. His guests gone, he talked to me of Bakst, a Russian with a lisp, whom I had once met at Mme de Pourtalès’.^{12} Bakst, a great decorator, illustrator, who made the costumes and designed the sets of the Ballets Russes. I was shown his colored sketches, some of them at least (for The Blue God, Narcissus, The Afternoon of a Faun, Sheherazade, etc.); Nijinsky had danced in them. When the history of the pre-war years gets written, these documents will certainly be mentioned for they reveal some queer, pathological states of mind, of those that herald revolutions.

7 February 1918

Lunch at Lucas’s with Mrs. Wharton, the Princesse Lucien Murat, the Baronne de Brimont, Berenson, Saint-André,^{13} Mac Lugan,^{14}, André Gide. André Gide sat next to me. A clean-shaven face, a balding head with the hair rather toward the back. Looks like a professor priest, a Protestant (which he is) with the stamp of austerity still upon him. Has been doing relief work since the outbreak of the war;^{15} is now in Normandy. He is married and does not look sorry for it. … On leaving the restaurant the Baronne de Brimont dragged me off to a small gallery on the rue du Faubourg-Saint-Honoré to a Picasso, Matisse, and Art Nègre exhibit. … Berenson did not appear to set great store by the new painters…

9 February 1918

Dined at Mrs. Wharton’s with Berenson, the architect Naville, Saint-André and Charpentier.^{16} Berenson remarked that André Gide’s face showed a strong resemblance to George Eliot’s. Naville who was a schoolmate of Gide at the École Alsacienne said that they had nicknamed him le crispatif there (neither crispé nor crispant, the invented crispatif may have meant that Gide’s tense face aggravated his classmates) because he talked with clenched teeth. Mrs. Wharton finds the author of The Immoralist shy. As I mentioned the austere cast of Gide’s face, Berenson added “a woman’s austerity, not a man’s.”

22 March 1918

Had lunch yesterday at Lucas’s with Mrs. Wharton, André Gide, Berenson. Gide told me of his strict puritanical upbringing and of his individualistic reaction afterwards. After a while, however, he recognized that it is the Gospel that redeems the individual. “For whosoever will save his life shall lose it.” The exaltation of art, he said, is the abnegation of art. … Gide mentioned Alfred Jarry’s Ubu Roi (‘King Ubu’)^{17} – a ferocious Rabelaisian eccentricity.

11 October 1923

Just back from the hotel Beau Site on the rue de Presbourg where I was Berenson’s guest for lunch. Mrs. Wharton talked to us about (the Abbey of) Pontigny^{18} where she spent a few hours. Silence is the rule there. Gide who had been informed that Paul Valéry was to be invited at the abbey objected to it on the grounds that Valéry was sure to talk too much. So, as I said, one does not talk there. Speaking of which, Berenson remarked: “I dislike collective cerebration.”

16 October 1923

Had lunch today at Walter Berry’s with the Princesse Lucien Murat, the Prince Bassiano,^{19} Jean Cocteau, Marie Laurencin,^{20} Mrs. Wharton. … Cocteau spoke quite highly of Paul Bourget. He has read him, admired La Geôle (‘the jail’). Bourget, he said, was made for the news item. … He told Walter Berry and myself that the future of literature lay in clarity, clarity enriched by all the past complexities. It will be a new classicism, without any rococo. …

24 July 1924

Mr. and Mrs. Henraux drove me down to Versailles yesterday where I had lunch at the Prince and Princesse Bassiano’s with Mrs. Wharton and Walter Berry. … Berenson has always felt the burden of his Jewish origin, Lucien Henraux told me. He also said that once on hearing his name spoken Mme Mühlfeld had crossed herself; she can’t stand the sight of him. Good grief!

References to Walter Berry

21 April 1915 (see Wharton)

13 August 1915 (ibid.)

25 June 1917

Lunch at Mme Cocteau’s, with the Princesse Lucien Murat, the Comtesse de Chevigné, Jean Cocteau, and Walter Berry. Mme de Chevigné is the great-granddaughter of Laure de Noves, the great-niece of the Marquis de Sade.^{21} She said that the marquis worshiped his wife; all his letters to her are love letters.

28 February 1918

Dined at the hotel Ritz last night as a guest of the Princess Soutzo.^{22} With Marcel Proust we talked about hawthorns. … Berry told me that Berenson is quite unhappy at not being more successful with women. And speaking of women, he also said that the Ritz had become a brothel at the onset of the war. Berry added that he could not understand why the act of copulation was considered a crime. It is as natural as ingesting food, he said. I quite like Proust.

17 April 1918

Had lunch at Walter Berry’s with the Comtesse Joachim Murat and Jean Cocteau. … Jean Cocteau was discoursing about the times in which we are living. …

27 November 1918

Dinner last night at Walter Berry’s with Jean Cocteau and the Duc de Guiche. Berry had been expecting several charming ladies who did not turn up. Jean Cocteau had quite a few stories to tell about art dealers; he mentioned names. … Cocteau also spoke at some length of Jacques-Emile Blanche,^{23} quite uncritically. Blanche has sent the proofs of his coming book on some painters to Berry for his archives. There is a preface by Marcel Proust. …

19 February 1921

Dined at the Princesse Marthe Bibesco’s with Walter Berry. … We talked about Abel Bonnard.^{24} Lack of success makes him hard on others, Berry said; he’s lazy.

17 October 1921

Dined last night at Walter Berry’s. Jean Cocteau joined us during the meal. I spoke to him of his book on (or against) Barrès.^{25} He has considered Barrès as one would “a pelota court wall”: with espaliers of fruit and flowers on the reverse side. He calls Anna de Noailles:“Sainte Opportune.”

16 October 1923 (see Wharton)

24 July 1924 (ibid.)

11 June 1927

Hurried to the Beaumonts’^{26} who had assembled a very brilliant assembly to view a lengthy German film, Metropolis. Bizarre, crazy, etc. Factories, offices... A tower of Babel with strange men, women, machines. Clusters of people, angry crowds, ascendings, catastrophes, a flood, dances, a cathedral and underground passages; from time to time men’s lips coming near women’s lips. To be short, a series of nightmares. “Quite teutonic,” was the general verdict. Technically, of course, it has required some tremendous labor. Without forgetting the music, often quite bewildering -snatches of Wagner’s Tetralogy to accompany ultra-modern, ultra-futurist sequences. Left around 1 a.m. Walter Berry took pity on me and kindly drove me home.

References to Bernard Berenson

21 November 1917

Berenson was telling me of Walter Pater’s enormous influence on himself and his generation. Pater, a man of great propriety, touring Italy with two deadly boring sisters. With a face like that of “a man of the earth.” A humanist, a Hellenist. Oscar Wilde was influenced by him though he never became a disciple. Wilde is “less refined.” Marius the Epicurean is about the death throes of Paganism and the birth of Christianity. Pater was a “fellow” at a college. He was in his sixties when he died. His influence was greater in the United States than in England. He wrote some exquisite pages on Watteau,^{27} Winckelmann^{28} and others. There is also something on Plato. He gives you the straighforward impression, the aesthetic one. Berenson knew Oscar Wilde well. He was an Irishman with something of the trickster about him, of the exhibitionist. He was a snob. At the Café Royal in London where they had lunch together, he was saying to Berenson: “I am presently to show myself at the Countess’; it’s tiresome, but it’s necessary.” Wilde is a mediocre author. He was on the verge of becoming a decent playwright when he died. Berenson came across portraits of him in books on homosexuality published in Berlin. In Germany there are many “Spartan homosexuals.”

5 January 1918

Lunch at Mr. Berenson’s today with Miss Barney^{29} and a Pole, Mr. Milosz,^{30} who has the shaven face of an actor, of a Dominican monk. I’m using the epithet on purpose, for Berenson told me afterwards that he intends to become a Dominican. He has the typical Dominican physiognomy. … Berenson believes that Goethe underwent some kind of ossification on his return from Italy and that he became pompous (‘pompier’) in a way, which he wasn’t before he went. The ladies around him fell into bigotry and so, at eighty, we see him in a tavern raising his glass to Paganism. … Berenson thinks that the Reims cathedral is much more beautiful with the wounds inflicted to it by the war. … And I left these charming guests to ride the subway home to my lodgings.

7 February 1918 (see Wharton)

9 February 1918 (ibid.)

28 February 1918 (see Berry)

22 March 1918 (see Wharton)

11 October 1923 (ibid.)

24 July 1924 (ibid.)

Notes

1 R.W.B. Lewis, Edith Wharton. A Biography (New York: Harper & Row, 1975); all references to it will appear parenthetically in the text of this article.

2 it is also the site of many government offices and of the French Assemblée Nationale. The Prime Minister’s Hôtel Matignon is located at No. 57 and 58 of the rue de Varenne. No. 58 was Wharton’s own address in 1907 and 1908.

3 Paris: Le Mercure de France.

4 La Vie d’une Amitié, 3 vol. (Paris: Plon, 1951-57).

5 Paris: Grasset, 1970.

6 At 53, rue de Varenne which she had leased in the spring of 1909 and where she was to live until she moved to Pavillon Colombe at Saint Brice-sous-Forêt north of Paris in 1920.

7 Abbé Henri Brémond (1865-1935): an ex-Jesuit, author of a voluminous (yet unfinished) Literary History of the Religious Feeling in France from the Wars of Religion to the Present Day. Elected to the French Academy in 1923.

8 A strange statement, considering the full title of James’s book, William Wetmore Story and His Friends: From Letters, Diaries and Recollections, which appeared in two volumes in 1903. Story (1819-1895), a native Bostonian and a lawyer by profession before settling in Rome in 1856, was famous for his statue of Cleopatra which was exhibited at the London exposition of 1862. James found his statues “not altogether unsuggestive of Mrs Jarley’s waxworks” and Mrs Henry Adams, on visiting his studio, regretted that “he spoiled so many… nice blocks of white marble” (quoted in Leon Edel, Henry James: A Life (New York: Harper and Row 1985), 140-41).

9 The Abbé mentions him one more time in his Journal in the entry for 28 May 1924 when he records the presence at a dinner he was attending of the writer Colette “clad in a printed fabric by Rodier with some red in it, her arms bare and strong.” Might Rodier have been a fabric designer or a couturier, in some way related to the current knitwear label Rodier?

10 Nancy Astor, later Lady Astor (1879-1964), the second wife of William Waldorf Astor whose father had become a British citizen in 1890 and was to be made the first Viscount Astor in 1917. In 1919, she was to become the first woman member of the British Parliament. She was the brilliant mistress of Cliveden, her country house in England, where Wharton was a frequent visitor.

11 Walter Berry’s residence at the time was 14, rue Saint-Guillaume, a street perpendicular to the Boulevard Saint-Germain. He was to take up the lease of Wharton’s apartment at 53, rue de Varenne when she moved to St. Brice in 1920.

12 Perhaps the Comtesse Mélanie de Pourtalès whose salon was located on the rue Tronchet, outside the Faubourg. She was the wife of a Lutheran banker of Swiss origin and had been a lady-in-waiting of the empress Eugénie, the wife of Napoleon III. Wharton had met her in the early 1880s at Cannes on the French Riviera when she was a young girl, and then later in Paris during her bouts of socializing in the winter of 1908 (Lewis 195). Mme de Pourtalès had died the previous year. It may also be the Comtesse Jacqueline who frequently entertained the Abbé or the Comtesse Hélène.

13 R. W. B. Lewis mentions Alfred de Saint-André as a frequent guest or co-guest of Wharton’s who became a “regular” like the Bourgets, the du Boses or the Comtesse de Fitz-James. He describes him as “a man of no visible achievements” but “a great gourmet and a connoisseur of out-of-the-way restaurants” (Lewis 196-97).

14 Mac Lugan must be Eric Maclagan, another “regular” of Wharton’s set. The head of the British Ministry of Information bureau in Paris, he was attached to the British Peace Delegation in 1919, and later became the director of the Victoria and Albert Museum in London (Lewis 408 & 476). He was another fine gourmet and discoverer of little restaurants, and also “a skilled raconteur of bawdy stories” (says Lewis), two qualities one would not basically expect from the son of an archbishop (his father, William Maclagan, had been the Archbishop of York).

15 Gide and Wharton had met in connection with the American Hostels for Refugees which Wharton had established at the very onset of the war. Gide became a member of the Franco-American General Committee of which Wharton was the chairperson. They eventually became friends and the two of them had spent several days in the company of each other in and around Hyères (on the Mediterranean in Provence) in the late fall of 1915 (Lewis 372). They saw each other regularly during the twenties.

16 One would think of the composer of the opera Louise (1900), Gustave Charpentier (1860-1957), but he may also be the publisher referred to in a later entry of the Journal (for 19 February 1925) to whom the novelist Alphonse Daudet would send letters of recommendation for young aspiring writers, adding a secret code for his correspondent not to act upon them.

17 Alfred Jarry (1873-1907) wrote a number of novels and comedies in his short paroxystic life, quite a few around the central farcical character of the tyrant Ubu (an avatar of one of his secondary school teachers). The opening of Ubu Roi on 10 Dec. 1896 had been a stormy affair.

18 An old Cistercian abbey on the northern confines of Burgundy founded in 1114. The remaining buildings served as a kind of retreat for writers and intellectuals after the war under the aegis of Paul Desjardins (1859-1940). Lewis mentions one such visit of Wharton’s during a motor trip to Nohant (George Sand’s château in the central region of Berry) with Walter Berry in the early twenties. She found herself at lunch in the company of Gide, du Bos, André Maurois and Lytton Strachey talking about the art of translation (Lewis 455-56).

19 The Prince & Princess Bassiano had a literary salon at Versailles where Wharton was a frequent guest after she moved to St. Brice. Lewis writes (439) that she reencountered André Gide there and that it was at the Bassianos’ that she first met Paul Valéry who then came periodically to visit her, both at St. Brice and at her Château Sainte-Claire in Hyères on the Mediterranean.

20 The painter (1885-1956) who (almost) never painted the noses of her models. She frequented the painters of the Bateau Lavoir in Paris (cf. her picture “Apollinaire and his friends”) and was the poet Apollinaire’s companion for a while.

21 The Comtesse de Chevigné is assumed to be the model for Proust’s Duchesse de Guermantes, but so is also the Comtesse Greffulhe.

22 The Princess Hélène Soutzo, née Chrissoveloni , was to marry the writer Paul Morand in 1927.

23 Jacques-Emile Blanche (1861-1942) was not only the fashionable portrait painter we know from his portraits of Henry James, Proust, Cocteau, Gide and others, he was also an art critic and a memorialist of his time (Propos dans l’Atelier; Cahiers d’un Artiste). Wharton was a frequent guest at his villa in Passy (Paris) and his manor house at Offranville in Normandy.

24 A writer and journalist (1883-1968) who turned to politics and became Minister of Education during the Vichy régime. He was sentenced to death after the war but had the sentence reduced to ten years of banishment. He’s not to be confused with the painter Pierre Bonnard.

25 Maurice Barrès (1862-1923), a writer and politician, a staunch supporter of nationalistic views and of the war against Germany. The Abbé Mugnier’s Journal mentions him as many as 81 times from 1891 on. He had a short affair with Anna de Noailles and there was a nasty scandal over the suicide of his nephew Charles Demange in 1909.

26 The Comte & Comtesse Etienne de Beaumont in their town house of the rue Duroc and the Vicomte & Vicomtesse Charles de Noailles in theirs at 3, Place des États-Unis (which was once Wharton’s brother Harry’s residence and on occasions her own) were the patrons of the artists during the “Années Folles” (the “Roaring Twenties”) in France. They participated in many of the Surrealists’s pranks, sponsored many of Jean Cocteau’s films and plays. The Abbé mentions the scandal around Cocteau’s The Blood of a Poet at the Noailles’, the viewing of which was hastily cancelled on the very night of its scheduled projection with the Abbé’s attendance (on 17 November 1930, according to the Abbé’s diary).

27 The French painter (1684-1721) of the Fêtes Galantes (The Embarkment for Cythera, Gilles, etc.).

28 Johann Joachim Winckelmann, the German archaeologist and art historian (1707-1768), the advocate of a return to the simplicity of Greek art which opened the way to Neo-Classicism in literature and the arts.

29 Natalie Barney was “the highly cultivated Ohio-born woman who presided over a cult of Lesbos and a high-powered literary salon in rue Jacob”, to steal Lewis’s own definition (407). It seems that Wharton made a point of staying clear of her, although Lewis mentions a lunch in Walter Berry’s apartment “in the fall of 1926” where they were co-guests (Lewis 444). The Abbé Mugnier appears not to have shared Wharton’s reluctance and Berenson was one of her gentlemen friends.

30 Oscar Vladislav de Lubics Milosz (1877-1939) was of Lithuanian origin like Berenson and lived in Paris where he wrote poetry, plays and novels.

Université des Sciences Humaines, Strasbourg, France.